# Privacy Guarantees in Posterior Sampling under Contamination

### v2

@@ 2.1 Differential Privacy in Bayesian Inference
Hereafter, we will assume the dataset to be precontaminated and denote $\bm{X}$ as the result of the contamination. Thus the hypothetical neighbouring dataset $\bm{Z}$ is a neighbour to the already contaminated dataset $\bm{X}$ and thus $\bm{Z}$, obtained from arbitrarily changing one entry of $\bm{X}$, contains the same contamination as $\bm{X}$ (possibly except for one entry).
2.3 Comparing frequentist and Bayesian approaches in DP
Before we introduce the main result, we present a motivating example for seeking a differential privacy bound that does not require assumptions of bounded parameter or observation space. The most common frequentist approach to retaining privacy is to add noise directly to the output (Dwork et al., 2006, 2014). Clearly such an approach will (like our contamination methodology) lose statistical efficiency. Within this simple example, we can directly compare the loss of efficiency that results from insisting on particular levels of DP.
The level of noise added by frequentist approaches usually grows with the size of the bounded space. Thus, this example shows how in the finite sample regime, the Bayesian approach can yield better statistical efficiency than the frequentist approaches especially with small datasets, even though asymptotically no Bayesian mechanism can beat the optimal frequentist estimator, as recognised in Wang, Fienberg and Smola (2015), due to Bernstein-von Mises theorem (Van der Vaart, 2000).
Consider the following simple linear regression model with no covariates:
$x_{i}=\theta+e_{i}$
where the true value $\theta^{*}$ is assumed to be $30$, and $e_{i}\sim\mathcal{N}_{[l,u]}(0,\sigma^{2})$ are i.i.d. truncated normal distributions supported on interval $[l,u]$, with standard deviation $\sigma=8$.
Let $\bm{X}:=\{x_{1},\dots,x_{n}\}$ be the set of observations, for given contamination rate $p_{n}:=\exp(-n^{1/12})$, the dataset is contaminated by replacing with probability $p_{n}$ each observation using $\tilde{x}_{i}\sim T_{\nu}(\theta^{*},\sigma)$ supported on $[l,u]$. The density of the scaled and truncated t-distribution is given by
Then we empirically derive $\hat{\epsilon}$ and $\hat{\delta}$ with respect to the dataset by using the approach we will describe later in Section 4. Finally, we produce the following three different estimations of $\theta$ with respect to the same privacy level $(\hat{\epsilon},\hat{\delta})$:
  1. 1.
    Bayesian: draw $\hat{\theta}_{B}\sim\pi_{n}(\cdot|\bm{X})$;
  2. 2.
    Frequentist (Laplace noise; Dwork et al. (2006)):
    $\hat{\theta}_{L}:=\frac{1}{n}\sum_{i=1}^{n}x_{i}+\text{Laplace}(\tfrac{u-l}{n\hat{\epsilon}});$ (4)
  3. 3.
    Frequentist (Guassian noise; Dwork et al. (2014)):
    $\hat{\theta}_{N}:=\frac{1}{n}\sum_{i=1}^{n}x_{i}+\text{Normal}\left(0,2\log\tfrac{1.25}{\hat{\delta}}\left(\tfrac{u-l}{n\hat{\epsilon}}\right)^{2}\right).$ (5)
Then by definition, $\hat{\theta}_{B}$ and $\hat{\theta}_{N}$ are $(\hat{\epsilon},\hat{\delta})$-DP and $\hat{\theta}_{L}$ is $(\hat{\epsilon},0)$-DP.
Since the implementation of Laplace or Gaussian noise requires a boundedness assumption observations $x_{i}\in[l,u]$, $\forall i$. To demonstrate the effect of changing the interval size, the simulation is conducted on two different support intervals $[l,u]=[0,300]$ and $[l,u]=[0,600]$. For both support sizes, the simulation is repeated 50 times, with respect to the same dataset $\bm{X}$, and the empirical variance of the estimators is computed. The results are summarized in Tables 0(a) and 0(b) respectively.
Since the frequentist approaches are unbiased, but the Bayesian approach is biased, we use the Mean Squared Error (MSE) as a measure of statistical efficiency. Since both frequentist estimators are unbiased, the MSE is dominated by the variance of the injected noise. Noticeably, both (frequentist) estimators roughly quadrupled in variance as the interval size doubled, which matches (4) and (5). On the other hand, we notice that the Bayesian estimator is mostly unaffected, both in empirical estimates of $\hat{\epsilon}$, $\hat{\delta}$ and in the MSE, when the interval size doubled.
In addition, the contamination probabilities are roughly $0.23$ and $0.17$ when $n=100$ and $1000$ respectively. Nonetheless, in these cases, the Bayesian approach still offers a better trade-off between statistical efficiency and privacy since the estimator has far better MSE compared with the two frequentist approaches.
This shows that while the frequentist’s estimator can achieve the optimal variance when the Cramer-Rao lower bound is attained in the asymptotic regime, a conservative bound assumption can lead to unnecessarily large statistical noise, especially when the dataset is moderately small. In contrast, the Bayesian approach gives a more robust estimator, in the sense that the level of differential privacy and estimator variance it achieves is independent of the size of the underlying space.
Table 1: Mean Square Error Comparison for Frequentist and Bayesian Estimators
$n$ $\hat{\epsilon}$ $\hat{\delta}$ MSEL MSEG MSEB
100 0.34 0.61 164.67 133.52 4.48
1000 0.23 0.22 3.34 6.54 0.33
5000 0.19 0.027 0.24 0.95 0.07
10000 0.16 0.003 0.07 0.47 0.03
(a) $l=0$, $u=300$
$n$ $\hat{\epsilon}$ $\hat{\delta}$ MSEL MSEG MSEB
100 0.34 0.61 658.08 537.81 4.30
1000 0.23 0.22 12.45 26.15 0.35
5000 0.19 0.03 0.87 3.73 0.07
10000 0.17 0.003 0.27 1.71 0.03
(b) $l=0$, $u=600$
@@ 2.4 Posterior Decomposition
$d(\bm{x},\theta):=\frac{k_{p}(\bm{x},\theta)}{k_{p}(\bm{x},\theta^{*})}$
where $\theta^{*}$ is the true parameter. If $X$ is a dataset with $n$ data points $x_{i},i=1,\dots,n$, then for $A\subset\Omega$, denote
$m_{n}(A,\bm{X})=\int_{A}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\theta)\mathtt{d}\theta.$
Now, we can write the posterior distribution function $\pi_{n}(\cdot|\bm{X})$ in terms of $m_{n}(\cdot,\bm{X})$,
$\mathbb{P}_{\pi_{n}}(A|\bm{X})=\frac{\int_{A}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}{\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}=\frac{m_{n}(A,\bm{X})}{m_{n}(\Omega,\bm{X})}.$ (6)
Let $A_{n}:=\{\theta\in\Omega:h(\theta,\theta^{*})\leq\phi_{n}\}$, defined by a positive sequence $\{\phi_{n},n\in{\mathbb{N}}\}\rightarrow 0$ and $h$ is the Hellinger distance defined (later) in Definition 3. $A_{n}$ may be defined using other distance metrics such that $\mathbb{P}_{\pi_{n}}(A_{n}^{c})\rightarrow 0$. We will prove our main theorems for Hellinger distance but use $L_{\infty}$ distance in the simulation. Then
$\mathbb{P}_{\pi_{n}}(S|\bm{X})=\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{X})+\mathbb{P}_{\pi_{n}}(S\cap A_{n}^{c}|\bm{X})$
@@
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\eta_{n}\left[\eta_{n}+\frac{m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\right]\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}$ (7)
@@
Now if we can find the upper bounds for $m_{n}(A_{n}^{c},\bm{X})$, $m_{n}(A_{n}^{c},\bm{Z})$ and $\eta_{n}$, and the lower bound for $m_{n}(\Omega,\bm{X})$ to upper bound the above expression, then we can treat (the upper bounds of) $\eta_{n}(\eta_{n}+\tfrac{m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})})$ as $e^{\epsilon}$ and $\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}$ as $\delta$.
@@
$\mathbb{P}_{\pi_{n}}(S\cap A_{n}^{c}|\bm{X})\leq\mathbb{P}_{\pi_{n}}(A_{n}^{c}|\bm{X})=\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}$
@@
$\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{X})\leq$ $\,\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{\int_{A_{n}\cap S}\prod_{i=1}^{n}d(\bm{z}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}{\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}$
$\leq$ $\,\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}{\int_{\Omega}\prod_{i=1}^{n}d(\bm{z}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}\mathbb{P}_{\pi_{n}}(A_{n}\cap S|\bm{Z})$
$\leq$ $\,\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{m_{n}(A_{n},\bm{Z})+m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\mathbb{P}_{\pi_{n}}(A_{n}\cap S|\bm{Z})$ (9)
3 Generic Results on DP Guarantees under Bayesian Sampling
This section will be devoted to presenting our main results and discussing the required assumptions.
3.1 Preliminaries
When the parameter space is non-compact, we need some control over the space complexity of density functions to compute the tail bound for the posterior distribution, which will act as an upper bound of $\delta$. In empirical processes, covering numbers and bracketing numbers are often used to quantify the size and complexity of a metric space.
@@ 3.1 Preliminaries
Let $p,q$ be two density functions defined on a measurable space $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$, then the (squared) Hellinger distance between $p$ and $q$ is defined as
@@
When the context is clear, we will use $h(\theta_{1},\theta_{2})$ to denote the distance between two contaminated density functions $h(k_{p}(\cdot;\theta_{1}),k_{p}(\cdot;\theta_{2}))$, and replace the function space using the parameter space $\Omega$.
@@
A set of function pairs $\{[L_{j},U_{j}],j=1,\dots,N\}\subset\Xi\times\Xi$ is a $r$-Hellinger bracketing of the space $\mathcal{P}$, if
$h(L_{j},U_{j})\leq\epsilon\text{ for all }j\in\{1,\dots,N\}$
@@
An $\epsilon$-bracket is a pair $[L_{j},U_{j}]$ with $h(L_{j},U_{j})\leq r$.
@@
The Hellinger distance is a metric on density functions, but it also generalizes easily to a metric on the space of non-negative integrable functions for the purpose of the above definition.
@@
The $\epsilon$-Hellinger bracketing number is size of the smallest $\epsilon$-bracketing, defined as
$N_{[]}(\epsilon,\mathcal{P}):=\min\{|B|:B\text{ is an $\epsilon$-Hellinger bracketing of $\mathcal{P}$}\}.$
The $\bm{\epsilon}$-Hellinger bracketing entropy is the log of the bracketing number
$H_{[]}(\epsilon,\mathcal{P}):=\log N_{[]}(\epsilon,\mathcal{P}).$
3.2 Main Results
@@
Assumption 1 measures the space complexity of the density functions $f(\cdot;\theta)$. Note that $H_{[]}(u,\mathcal{F}(\Omega_{n}))\leq c^{2}n\phi_{n}^{2},\forall u\in(\phi_{n}^{2}/2^{10},\phi_{n})$ implies (10), then this condition can be verified easily with the following lemma.
@@
Let $\Omega_{1}\subset\Omega_{2}\subset\dots\subset\Omega$ be a sequence of subsets such that $\Omega_{n}\rightarrow\Omega$ and $\Omega$ is a convex subset of $\mathbb{R}^{d}$. Let $R_{n}$ denote the radius of the space $\Omega_{n}$ which grows at most polynomial order with $n$. If the set of density functions $\mathcal{F}(\Omega_{n})$ is locally Lipschitz in the sense that
$\left|\sqrt{f_{\theta}(\bm{x})}-\sqrt{f_{\vartheta}(\bm{x})}\right|\leq M^{(n)}_{\vartheta,u}(\bm{x})\|\theta-\vartheta\|_{\infty},$
@@
$M_{u}^{(n)}:=\sup_{\vartheta\in\Omega_{n}}\|M^{(n)}_{\vartheta,u}\|_{2}$
which is a constant depending on $R_{n}$. If there exists a constant $q\in(0,1)$, such that $\forall u\in(0,\phi_{0})$ for some constant $\phi_{0}$,
@@
then there exists positive sequence $\phi_{0}>\phi_{n}\rightarrow 0$ with $\phi_{n}>n^{-\tfrac{1-q}{4}}$ such that (10) holds.
@@
We will see later in the examples that $M_{u}^{(n)}$ grows with $u$, so usually it suffices to just check (11) for only $u=\phi_{0}$.
@@
The positive sequence $\phi_{n}\rightarrow 0$ satisfies satisfies
@@
for some constants $\alpha\in(0,1]$, $c_{3}>0$, where $t_{n}:=c_{1}\phi_{n}^{2}/8$ and $\tfrac{nt_{n}}{\log n}>\max\left\{\tfrac{1}{\alpha},\tfrac{c_{1}}{8c_{2}}\right\}$, $\forall n>0$, $c_{2}$ is a constant in Lemma 9.
@@
Assumptions 2 and 3 control the prior concentration on the center and the tails respectively. The first three assumptions are needed in order to control the $\delta$ term in (3). Note that Assumptions 1-3 feature the same symbol $\phi_{n}$ for the sequence. In the proof for the main result, we need the same sequence $\phi_{n}$ to satisfy these three assumptions simultaneously.
@@
Denote that supremum of the directional derivative of $f(\bm{x};\theta)$ in $\theta$ at $(\bm{x},\theta)$ by ~$\tilde{\nabla}_{\theta}f(\bm{x};\theta)$,
~$\tilde{\nabla}_{\theta}f(\bm{x};\theta)=\sup_{\|\vec{u}\|_{2}=1}\lim_{\xi\downarrow 0}\frac{f(\bm{x};\theta+\xi\vec{u})-f(\bm{x};\theta)}{\xi}.$
@@
All the directional derivatives of the density function $f(\bm{x};\theta)$ in $\theta$ exist and are bounded, and the contamination density $g$ has a tail no lighter than ~$\tilde{\nabla}_{\theta}f(\bm{x};\theta)$ in every direction, i.e., for any fixed $\theta\in\Omega$,
$\sup_{\bm{x}\in\mathcal{X}}\left\|\frac{\tilde{\nabla}_{\theta}f(\bm{x};\theta)}{g(\bm{x})}\right\|_{2}<\infty,$
@@
We assume a weaker condition than total differentiability to accommodate the use of absolute value ||$\cdot$|| ⋅ | in density functions, e.g., the Laplace distribution.
When $g$ is continuous and has full support over the support of ~$\tilde{\nabla}_{\theta}f$, then the assumption holds trivially whenever $g$ has a suitably heavy tail.
@@
$\inf_{\|\theta-\theta^{*}\|_{2}>\psi}h(f(\cdot;\theta),f(\cdot;\theta^{*}))>0$ (14)
@@
Assumption 5 is similar to one of the conditions in Thm 9.13 Wasserman (2004), but stronger, since we stated it in Hellinger distance rather than in KL divergence. The condition directly implies the identifiability of the model. See also Condition 2 in the lecture notes Wasserman (2020).
@@
If Assumption 5 holds, then for any sequence $\phi_{n}\rightarrow 0$, and decreasing sequence $1>p_{n}\rightarrow 0$, there exists sequence $\psi_{n}\rightarrow 0$ such that
@@
Theorem 5 ($(\epsilon,\delta)$-Privacy).
Suppose that there exists a decreasing sequence $\phi_{n}>0$ that satisfies the conditions in Assumptions 1-3 and that Assumptions 4, 5 also hold. For $p_{0}\in(0,1)$ fixed, there exists bounded positive constants (or sequence of constants) $\{T_{n}\},c_{1},c_{2},\dots,C_{1},C_{2},\dots<\infty$ such that $\forall S\in\mathcal{B}(\Omega)$ and any sequence $p_{0}\geq p_{n}\downarrow 0$,
$\mathbb{P}_{\theta^{*}}\Bigl{(}\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq e^{\epsilon_{n}}\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}\Bigr{)}\geq\begin{dcases}1-C_{1}e^{-nc_{5}\phi^{2}_{n}},&\text{for }nc_{2}\phi_{n}^{2}\geq\log 2,\\ 1-(C_{2}+C_{3}(\log\phi_{n})^{2})e^{-nc_{5}\phi^{2}_{n}},&\text{otherwise}\end{dcases}$ (16)

### v3

@@ 2.1 Differential Privacy in Bayesian Inference
Hereafter, we will assume the dataset to be precontaminated and denote $\bm{X}$ as the result of the contamination. Thus, the hypothetical neighbouring dataset $\bm{Z}$ obtained by changing one entry of the already contaminated dataset $\bm{X}$ carries the same contamination as $\bm{X}$ except possibly for one entry.
@@ 2.3 Posterior Decomposition
$d(\bm{x};\theta):=\frac{k_{p}(\bm{x};\theta)}{k_{p}(\bm{x};\theta^{*})},$
where $\theta^{*}$ is the true parameter. If $\bm{X}$ is a dataset with $n$ data points $x_{i},i=1,\dots,n$, then for $A\subset\Omega$, denote
$m_{n}(A,\bm{X})=\int_{A}\prod_{i=1}^{n}d(\bm{x}_{i};\theta)\pi_{0}(\theta)\mathtt{d}\theta,$
where the dependence on $\theta^{*}$ is omitted, since $\theta^{*}$ is considered constant in the model. Now, we can write the posterior distribution function $\pi_{n}(\cdot|\bm{X})$ in terms of $m_{n}(\cdot,\bm{X})$,
$\mathbb{P}_{\pi_{n}}(A|\bm{X})=\frac{\int_{A}\prod_{i=1}^{n}d(\bm{x}_{i};\theta)\pi_{0}(\mathtt{d}\theta)}{\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i};\theta)\pi_{0}(\mathtt{d}\theta)}=\frac{m_{n}(A,\bm{X})}{m_{n}(\Omega,\bm{X})}.$ (4)
Let $A_{n}:=\{\theta\in\Omega:h(\theta,\theta^{*})\leq\phi_{n}\}$, for a positive sequence $\phi_{n}\rightarrow 0$, where $h$ denotes the Hellinger distance which will be defined in Definition 3. $A_{n}$ may be alternatively defined using other distance metrics such that $\mathbb{P}_{\pi_{n}}(A_{n}^{c}|\bm{X})\rightarrow 0$. We will prove our main theorems for Hellinger distance, but use $L_{\infty}$ distance in the simulation. Then
$\mathbb{P}_{\pi_{n}}(S|\bm{X})=\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{X})+\mathbb{P}_{\pi_{n}}(S\cap A_{n}^{c}|\bm{X}),$
@@
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\eta_{n}\left[\eta_{n}+\frac{m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\right]\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})},$ (5)
@@
If we can find upper bounds for $m_{n}(A_{n}^{c},\bm{X})$, $m_{n}(A_{n}^{c},\bm{Z})$ and $\eta_{n}$, and lower bound for $m_{n}(\Omega,\bm{X})$, then we can treat (the upper bounds of) $\eta_{n}(\eta_{n}+\tfrac{m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})})$ as $e^{\epsilon}$ and $\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}$ as $\delta$.
@@
$\mathbb{P}_{\pi_{n}}(S\cap A_{n}^{c}|\bm{X})\leq\mathbb{P}_{\pi_{n}}(A_{n}^{c}|\bm{X})=\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})},$
@@
$\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{X})\leq$ $\,\sup_{\bm{x},\bm{z}\in\mathcal{X}}\left[\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\right]\frac{\int_{A_{n}\cap S}\prod_{i=1}^{n}d(\bm{z}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}{\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}$
$\leq$ $\,\sup_{\bm{x},\bm{z}\in\mathcal{X}}\left[\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\right]\frac{\int_{A_{n}\cap S}\prod_{i=1}^{n}d(\bm{z}_{i};\theta)\pi_{0}(\mathtt{d}\theta)}{\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i};\theta)\pi_{0}(\mathtt{d}\theta)}$
$\leq$ $\,\sup_{\bm{x},\bm{z}\in\mathcal{X}}\left[\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\right]\frac{\int_{\Omega}\prod_{i=1}^{n}d(\bm{z}_{i};\theta)\pi_{0}(\mathtt{d}\theta)}{\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i};\theta)\pi_{0}(\mathtt{d}\theta)}\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{Z})$ (7)
$\leq$ $\,\sup_{\bm{x},\bm{z}\in\mathcal{X}}\left[\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\right]\frac{m_{n}(A_{n},\bm{Z})+m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{Z})$
$\leq$ $\,\eta_{n}\left(\eta_{n}\mathbb{P}_{\pi_{n}}(A_{n}|\bm{X})+\frac{m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\right)\mathbb{P}_{\pi_{n}}(s\cap A_{n}|\bm{Z}),$
where the last step comes from drawing the supremum of the fraction $\tfrac{d(\bm{z};\theta)}{d(\bm{x};\theta)}$ out of the integral to transform $m_{n}(A_{n},\bm{Z})$ into $m_{n}(A_{n},\bm{X})$. ∎
2.4 Hellinger Distance and Bracketing Entropy
When the parameter space is non-compact, we need some control over the space complexity of density functions to compute the tail bound for the posterior distribution, which will act as an upper bound for $\delta$. In empirical processes, covering numbers and bracketing numbers are often used to quantify the size and complexity of a metric space.
@@ 2.4 Hellinger Distance and Bracketing Entropy
Let $p,q$ be two density functions defined on a measurable space $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$, then the squared Hellinger distance between $p$ and $q$ is defined as
@@
When the context is clear, we will use $h(\theta_{1},\theta_{2})$ to denote the distance between two contaminated density functions $h(k_{p}(\cdot;\theta_{1}),k_{p}(\cdot;\theta_{2}))$, and replace the function space with the parameter space $\Omega$.
@@
A set of function pairs $\{[L_{j},U_{j}],j=1,\dots,N\}\subset\Xi\times\Xi$ is a $r$-Hellinger bracketing of the space $\mathcal{P}$ if
$h(L_{j},U_{j})\leq r\text{ for all }j\in\{1,\dots,N\},$
@@
An $r$-bracket is a pair $L_{j},U_{j}]$ with $h(L_{j},U_{j})\leq r$.
@@
The Hellinger distance is a metric on density functions, but it also generalises easily to a metric on the space of non-negative integrable functions for the purpose of the above definition.
@@
The $r$-Hellinger bracketing number is the size of the smallest $r$-bracketing, defined as
$N_{[]}(r,\mathcal{P}):=\min\{|B|:B\text{ is an $r$-Hellinger bracketing of $\mathcal{P}$}\}.$
The $r$-Hellinger bracketing entropy is the log of the bracketing number
$H_{[]}(r,\mathcal{P}):=\log N_{[]}(r,\mathcal{P}).$
3 Main Results
This section is devoted to presenting our main results and discussing the required assumptions.
3.1 Main Assumptions
@@
Assumption 1 measures the space complexity of the density functions $f(\cdot;\theta)$. Note that $H_{[]}(u,\mathcal{F}(\Omega_{n}))\leq c^{2}n\phi_{n}^{2},\forall u\in(\phi_{n}^{2}/2^{10},\phi_{n})$ implies (8), thus this condition can be verified with the following lemma.
@@
Let $\Omega_{1}\subset\Omega_{2}\subset\dots\subset\Omega$ be a sequence of subsets such that $\Omega_{n}\rightarrow\Omega$ and $\Omega$ is a convex subset of $\mathbb{R}^{d}$. Let $R_{n}$ denote the radius of $\Omega_{n}$, which grows at most polynomially with $n$. Suppose that the set of density functions $\mathcal{F}(\Omega_{n})$ is locally Lipschitz in the sense that
$\left|\sqrt{f_{\theta}(\bm{x})}-\sqrt{f_{\vartheta}(\bm{x})}\right|\leq M^{(n)}_{\vartheta,u}(\bm{x})\|\theta-\vartheta\|_{\infty}$
@@
$M_{u}^{(n)}:=\sup_{\vartheta\in\Omega_{n}}\|M^{(n)}_{\vartheta,u}\|_{2},$
which is a constant depending on $R_{n}$. If there exists a constant $q\in(0,1)$ such that $\forall u\in(0,\phi_{0})$ for some constant $\phi_{0}$,
@@
then there exists a positive sequence $\phi_{0}>\phi_{n}\rightarrow 0$ with $\phi_{n}>n^{-\tfrac{1-q}{4}}$ such that (8) holds.
@@
Later, in the examples, we show that $M_{u}^{(n)}$ grows with $u$, so it usually suffices to check (9) for only $u=\phi_{0}$.
@@
The positive sequence $\phi_{n}\rightarrow 0$ satisfies
@@
for some constants $\alpha\in(0,1]$, $c_{3}>0$, where $t_{n}:=c_{1}\phi_{n}^{2}/8$, $\tfrac{nt_{n}}{\log n}>\max\left\{\tfrac{1}{\alpha},\tfrac{c_{1}}{8c_{2}}\right\}$, $\forall n>0$, and $c_{2}$ is a constant from Lemma 8.
@@
Assumptions 2 and 3 control the prior concentration around the centre and in the tails, respectively. The first three assumptions are needed to control the $\delta$ term in (3). Note that Assumptions 1-3 feature the same symbol $\phi_{n}$ for the sequence. In the proof of the main result, we require the same sequence $\phi_{n}$ to satisfy these three assumptions simultaneously.
@@
Denote the supremum of the directional derivative of $f(\bm{x};\theta)$ with respect to $\theta$ at $(\bm{x},\theta)$ by $\tilde{\nabla}_{\theta}f(\bm{x};\theta)$,
$\tilde{\nabla}_{\theta}f(\bm{x};\theta)=\sup_{\|\vec{u}\|_{2}=1}\lim_{\xi\downarrow 0}\frac{f(\bm{x};\theta+\xi\vec{u})-f(\bm{x};\theta)}{\xi}.$
@@
All the directional derivatives of the density function $f(\bm{x};\theta)$ with respect to $\theta$ exist and are bounded, and the contamination density $g$ has tails no lighter than $\tilde{\nabla}_{\theta}f(\bm{x};\theta)$ in any direction, i.e., for any fixed $\theta\in\Omega$,
$\sup_{\bm{x}\in\mathcal{X}}\left\|\frac{\tilde{\nabla}_{\theta}f(\bm{x};\theta)}{g(\bm{x})}\right\|_{2}<\infty.$
@@
We assume a weaker condition than total differentiability to accommodate the use of absolute value $|\cdot|$ in density functions, e.g., the Laplace distribution.
When $g$ is continuous and has full support over the support of $\tilde{\nabla}_{\theta}f$, the assumption holds trivially whenever $g$ has suitably heavy tails.
@@
$\inf_{\|\theta-\theta^{*}\|_{2}>\psi}h(f(\cdot;\theta),f(\cdot;\theta^{*}))>0.$ (12)
@@
Assumption 5 is similar to one of the conditions in Theorem 9.13 of Wasserman (2004), but stronger, since we stated it in terms of Hellinger distance rather than KL divergence. The condition directly implies the identifiability of the model. See also Condition 2 in the lecture notes Wasserman (2020).
The following Proposition helps quantify the rate of convergence for $\epsilon_{n}$ within a Hellinger neighbourhood of $\theta^{*}$ by employing Cauchy’s mean value theorem on the likelihood ratio $d(\bm{x};\theta)$.
@@
If Assumption 5 holds, then for any sequence $\phi_{n}\rightarrow 0$, and any decreasing sequence $1>p_{n}\rightarrow 0$, there exists a sequence $\psi_{n}\rightarrow 0$ such that
@@
3.2 DP under Bayesian Sampling
Theorem 4.
Suppose that there exists a decreasing sequence $\phi_{n}>0$ that satisfies the conditions in Assumptions 1-3, and suppose also that Assumptions 4, 5 hold. For a fixed $p_{0}\in(0,1)$, there exist positive constants (or a sequence of constants) $\{T_{n}\},c_{1},c_{2},\dots,C_{1},C_{2},\dots<\infty$ such that for any sequence $p_{0}\geq p_{n}\downarrow 0$,
$\mathbb{P}^{*}_{\theta^{*}}\Bigl(\forall S\in\mathcal{B}(\Omega),$ $\,\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq e^{\epsilon_{n}}\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}\Bigr)$ (14)
$\geq$ $\,\begin{dcases}1-C_{1}e^{-nc_{5}\phi^{2}_{n}},&\text{for }nc_{2}\phi_{n}^{2}\geq\log 2,\\ 1-(C_{2}+C_{3}(\log\phi_{n})^{2})e^{-nc_{5}\phi^{2}_{n}},&\text{otherwise}\end{dcases}$